What is the percent concentration (%,m/v) after 250 ML of an 8% (m/v) NaOH solution is diluted with 750 ML of water?
1
Expert's answer
2020-03-30T10:53:30-0400

"x\%(m/v)" solution contains x grams of solute in 100 ml. of solution.

Thus, 250 ml of "8\% (m/v) NaOH" solution contains "8\/100*250=20 gms." of NaOH.

On adding, 750 ml. of water to this solution, volume of the solution becomes "250+750=1000ml."

Thus, the new diluted solution contains 20 gms. of NaOH in 1000ml. of solution.

Thus, concentration in %(m/v) "=20/1000=2/100=2\%(m/v)" . (Answer)